Cyril invested part of 185 000 at 7% and the rest at 9. If the interest from the 9% investment in 1450 more than the interest earned at the 7% investment. How much did he invest at the 9% rate?
Answer by checkley77(12569)

.09X=.07(185,000-X)+1,450
.09X=12,950-.07X+1,450
.09X+.07X=12,950+1,450
.16X=14,400
X=14,400/.16
X=90,000 ANOUNT INVESTED @ 9%.
185,000-90,000=95,000 INVESTED @ 7%
PROOF:
.09*90,000=.07*95,000+1,450
8,100=6,650+1,450
8,100=8,100